• Prove that 7n−17n−1 is a multiple of 6 for all n∈N.
1
Expert's answer
2021-06-17T18:48:08-0400

Solution:

Let P(n) be "7^{n}-1" is a multiple of 6 for all n∈N.

For n=1

P(1): "7^{1}-1=7-1=6" is a multiple of 6.

So, P(1) is true.

Now,we assume that P(k) is true.

P(k): "7^{k}-1" is a multiple of 6

"\Rightarrow 7^k-1=6m" [for some constant m] ...(i)

Now, we show that P(k+1) is true.

P(k+1): "7^{k+1}-1" is a multiple of 6.

Take "7^{k+1}-1"

"=7^k7-1
\\=(6m+1)7-1\ \ \ [from (i)]
\\=42m+7-1
\\=42m+6
\\=6(7m+1)"

which is clearly a multiple of 6.

Hence, by the principle of mathematical induction, given statement P(n) is true for all n ∈N